OpenStudy (cameronlovesmorgan):

Cary is 9 years older than Dan. In 7 years, the sum of their ages will equal 93. Find both of their ages now.

OpenStudy (studygurl14):

Set up two equation based on the given info: C = cary's age D = dan's age D + 9 = C (C + 7) + (D + 7) = 93 Solve the system of equations